# Long-Sequence Recommendation Models Need Decoupled Embeddings

### v1

@@ Abstract
Lifelong user behavior sequences, comprising up to tens of thousands of history behaviors, are crucial for capturing user interests and predicting user responses in modern recommendation systems. A two-stage paradigm is typically adopted to handle these long sequences: a few relevant behaviors are first searched from the original long sequences via an attention mechanism in the first stage and then aggregated with the target item to construct a discriminative representation for prediction in the second stage. In this work, we identify and characterize, for the first time, a neglected deficiency in existing long-sequence recommendation models: a single set of embeddings struggles with learning both attention and representation, leading to interference between these two processes. Initial attempts to address this issue using linear projections—a technique borrowed from language processing—proved ineffective, shedding light on the unique challenges of recommendation models. To overcome this, we propose the Decoupled Attention and Representation Embeddings (DARE) model, where two distinct embedding tables are initialized and learned separately to fully decouple attention and representation. Extensive experiments and analysis demonstrate that DARE provides more accurate search of correlated behaviors and outperforms baselines with AUC gains up to 9‰ on public datasets and notable online system improvements. Furthermore, decoupling embedding spaces allows us to reduce the attention embedding dimension and accelerate the search procedure by 50% without significant performance impact, enabling more efficient, high-performance online serving.
@@ 1 Introduction
In online services, system response delays can severely disrupt the user experience, making effcient handling of long sequences within a limited time crucial. A general paradigm employs a two-stage process (Pi et al., 2020): search (a.k.a. General Search Unit) and sequence modeling (a.k.a. Exact Search Unit). This method relies on two core modules: the attention module [^1] In this paper, “attention” refers to attention scores—the softmax output that weights each behavior., which measures the target-behavior correlation, and the representation module, which generates a discriminative representation of behaviors. The search stage uses the attention module to retrieve the top-k relevant behaviors, constructing a shorter sub-sequence from the original long behavior sequence [^2] The search stage can also be “hard” selecting behaviors by category, but we focus on soft search based on learned correlations for better user interest modeling.. The sequence modeling stage then relies on both modules to predict user responses by aggregating behavior representations in the sub-sequence based on their attention, thus extracting a discriminative representation involving both behaviors and the target. Existing works widely adopt this effective paradigm (Pi et al., 2020; Chang et al., 2023; Si et al., 2024).
@@
Attention is critical in the long-sequence recommendation, as it not only models the importance of each behavior for sequence modeling but, more importantly, determines which behaviors are selected in the search stage. However, in most existing works, the attention and representation modules share the same embeddings despite serving distinct functions—one learning correlation scores, the other learning discriminative representations. Our analysis reveals that, unfortunately, gradients of these shared embeddings are dominated by representation learning during training, and more concerning, gradient directions from two modules tend to conflict with each other. As a result, attention fails to capture behavior importance accurately, causing key behaviors to be mistakenly filtered out during the search stage (as shown in Sec. 4.3). Furthermore, gradient conflicts also degrade the discriminability of the representations (as shown in Sec. 4.4).
@@
Inspired by the use of separate query, key (for attention), and value (for representation) projection matrices in the original self-attention mechanism (Vaswani et al., 2017), we experimented with attention- and representation-specific projections in recommendation models, aiming to resolve conflicts between these two modules. However, this approach did not yield positive results and led to over-confidence in attention (as shown in Sec. 2.3). Through insightful empirical analysis, we hypothesize that the failure is due to the significantly lower capacity (i.e., fewer parameters) of the projection matrices in recommendation models compared to those in natural language processing (NLP). This limitation is difficult to overcome, as it stems from the low embedding dimension imposed by interaction collapse theory (Guo et al., 2023).
@@
To address these issues, we propose the Decoupled Attention and Representation Embeddings (DARE) model, which completely decouples these two modules at the embedding level by using two independent embedding tables—one for attention and the other for representation. This decoupling allows us to fully optimize attention to capture correlation and representation to enhance discriminability. Furthermore, by separating the embeddings, we can accelerate the search stage by 50% by reducing the attention embedding dimension to half, with minimal impact on performance. On the public Taobao and Tmall long-sequence datasets, DARE outperforms the state-of-the-art TWIN model across all embedding dimensions, achieving AUC improvements of up to 9‰. Online evaluation on one of the world’s largest online advertising platforms achieves a 1.47% lift in GMV (Gross Merchandise Value). Our contribution can be summarized as follows:
  • We identify the issue of interference between attention and representation learning in existing long-sequence recommendation models and demonstrate that linear projections borrowed from NLP fail to decouple these two modules effectively.
  • We propose the DARE model, which uses module-specific embeddings to fully decouple attention and representation. Our comprehensive analysis shows that our model significantly improves attention accuracy and discriminability of representations.
@@ 2 An In-Depth Analysis into Attention and Representation
In this section, we first review the general formulation for long-sequence recommendation. Then, we analyze the training of shared embeddings, highlighting the domination and conflict of gradients from the attention and representation modules. Finally, we explore why straightforward approaches using module-specific projection matrices fail to address the issue.
@@ 2.2 Gradient Analysis of Domination and Conflict
Figure 3: The cosine angles of gradients from two modules.
The attention and representation modules have distinct goals: the former focuses on learning correlation scores for behaviors, while the latter focuses on learning discriminative (i.e., separable) representations in a high-dimensional space. However, current methods use a shared embedding for both tasks, which may prevent either from being fully achieved. To validate this assumption, we analyze the gradients from both modules on the shared embeddings.
@@
We empirically observe the gradients back propagated to the embeddings from the attention and representation modules. Comparing their norms, we find that gradients from the representation are five times larger, dominating those from attention, as demonstrated in Fig. 3. Then, we further explore whether they can be consistent with each other by analyzing gradient directions. Unfortunately, results show that in nearly two-thirds of cases, the cosine of their angles is negative, indicating the conflict between them, as shown in Fig. 3. In summary, the attention module and representation modules optimize the embedding table towards different directions with varying intensities during training, causing attention to lose correlation accuracy and representation to fail to fully fulfill its discriminability. Notably, due to domination, such influence is more severe to attention, as indicated by the poor learned correlation between categories in Sec. 4.3.
Finding 1. The embedding gradients are typically dominated by the representation module. Furthermore, gradients from the attention and representation modules tend to conflict.
@@
Figure 4: Illustration and evaluation for adopting linear projections. (a-b) The attention module in original TWIN and after adopting linear projections. (c) Performance of TWIN variants. Adopting linear projections causes an AUC drop of nearly 2% on Taobao.
@@ 2.3 Recommendation Models Call for More Powerful Decoupling Methods
Figure 5: The dispersed distribution of attention scores in TWIN w/ proj.
Linear projections would cause dispersed distribution of attention score.
To address such conflict, a straightforward approach is to use separate projections for attention and representation, mapping the original embeddings into two new decoupled spaces. This is adopted in the standard self-attention mechanism (Vaswani et al., 2017), which introduces query, key (for attention), and value projection matrices (for representation). Inspired by this, we propose a variant of TWIN that utilizes linear projections to decouple attention and representation modules, named TWIN w/ proj.. The comparison with the original TWIN structure is shown in Fig. 4a and 4b. Surprisingly, linear projection, which works well in NLP, loses efficacy in recommendation systems, leading to negative performance impact, as shown in Tab. 4c. An analysis of attention logits distribution in Fig. 5 shows that the logits of the TWIN w/ proj have a very dispersed distribution. For more analysis, refer to Sec. 4.3.
@@
The failure of introducing projection matrices makes us wonder why it works well in NLP but not in recommendation. One possible reason is that the relative capacity of projection matrices regarding the token numbers in NLP is usually large, e.g., with an embedding dimension of 4096 in LLaMA3.1 (Dubey et al., 2024), there are around 16 million parameters ($4096\times 4096=16,777,216$) in each projection matrix to map only 128,000 tokens in the vocabulary. To validate our hypothesis, we conduct a synthetic experiment in NLP using nanoGPT (Andrej, ) with the Shakespeare dataset. In particular, we decrease its embedding dimension from 128 to 2 and check the performance gap between two models with/without projection matrix. As shown in Fig. 6, we observe that when the matrix has enough capacity, i.e., the embedding dimension is larger than 16, projection leads to significantly less loss. However, when the matrix capacity is further reduced, the gap vanishes. Our experiment indicates that projection matrix only works with enough capacity.
@@
In contrast, due to the interaction collapse theory (Guo et al., 2023), the embedding dimension in recommendation is usually no larger than 200, leading to only up to $40000$ parameters for each matrix to map millions to billions of IDs. Therefore, the projection matrices in recommendation never get enough capacity, making them unable to decouple attention and representation.
Finding 2. The linear projection matrices fail to decouple attention and representation in recommendation models due to limited capacity.
@@ 3 DARE: Decoupled Attention and Representation Embeddings
In the above analysis, we find that using projection matrices is insufficient to decouple the attention and representation. To this end, we propose to decouple these two modules at the embedding level; that is, we employ two embedding tables, one for attention ($\bm{E}^{\text{Att}}$) and another for representation ($\bm{E}^{\text{Repr}}$). With gradient back propagated to different embedding tables, our method has the potential to fully resolve the gradient domination and conflict between these two modules. We introduce our model specifically in this section and verify its advantage by experiments in the next section.
@@ 3.1 Attention Embedding
Attention measures the correlation between history behaviors and the target (Zhou et al., 2018). Following the common practice, we use the scaled dot-product function (Vaswani et al., 2017). Mathematically, the $i$-th history behavior $i$ and target $t$, are embedded into $\bm{e}_{i}^{\text{Att}},\bm{v}_{t}^{\text{Att}}\sim\bm{E}^{\text{Att}}$, where $\bm{E}^{\text{Att}}$ is the attention embedding table. After retrieval $1,2,\dots,K=\mathrm{argsort}(\langle\bm{e}_{i},\bm{v}_{t}\rangle,i\in\left[1,N\right])$ their weight $w_{i}$ is formalized as:
@@
Figure 7: Architecture of the proposed DARE model. One embedding is responsible for attention, learning the correlation between the target and history behaviors, while another is responsible for representation, learning discriminative representations for prediction. Decoupling these two embeddings allows us to resolve the conflict between the two modules.
@@ 3.2 Representation Embedding
In the representation part, another embedding table $\bm{E}^{\text{Repr}}$ is used, where $i$ and $t$ is embedded into $\bm{e}_{i}^{\text{Repr}},\bm{v}_{t}^{\text{Repr}}\sim\bm{E}^{\text{Repr}}$. Most existing methods multiply the attention weight with the representation of each retrieved behavior and then concatenate it with the embedding of the target as the input of Multi-Layer Perceptron (MLP): $[\sum_{i}w_{i}\bm{e}_{i},\bm{v}_{t}]$. However, it has been proved that MLP struggle to effectively learn dot product or explicit interactions (Rendle et al., 2020; Zhai et al., 2023). To enhance the discriminability, following TIN (Zhou et al., 2024), we adopt the target-aware representation $\bm{e}_{i}^{\text{Repr}}\odot\bm{v}_{t}^{\text{Repr}}$, which abbreviate as TR in our following paper (refer to Sec. 4.4 for empirical evaluation of discriminability).
The overall structure of our model is shown in Fig. 7. Formally, user history $h$ is compressed into:
$\bm{h}=\sum_{i=1}^{K}w_{i}\cdot(\bm{e}_{i}^{\text{Repr}}\odot\bm{v}_{t}^{\text{Repr}}).$ (2)
@@ 3.3 Inference Acceleration
By decoupling the attention and representation embedding tables, the dimension of attention embeddings $\bm{E}^{\text{Att}}$ and the dimension of representation embeddings $\bm{E}^{\text{Repr}}$ have more flexibility. In particular, we can reduce $\bm{E}^{\text{Att}}$ while keeping $\bm{E}^{\text{Repr}}$ to accelerate the searching over the original long sequence whilst not affecting the model’s performance. Empirical experiments in Sec. 4.5 shows that our model have the potential to speed up searching by 50% with quite little influence on performance and even by 75% with an acceptable performance loss.
@@ 3.4 Discussion
Considering the superiority of decoupling the attention and representation embeddings, one may naturally raise an idea: we can further decouple the embeddings of history and target within the attention (and representation) module, i.e. forming a TWIN with 4 Embeddings method, or TWIN-4E in short, consisting of attention-history (named keys in NLP) $\bm{e}_{i}^{\text{Att}}\in\bm{E}^{\text{Att-h}}$, attention-target (named querys in NLP) $\bm{v}_{t}^{\text{Att}}\in\bm{E}^{\text{Att-t}}$, representation-history (named values in NLP) $\bm{e}_{i}^{\text{Repr}}\in\bm{E}^{\text{Repr-h}}$ and representation-target $\bm{v}_{t}^{\text{Repr}}\in\bm{E}^{\text{Repr-t}}$. The structure of TWIN-4E is shown in Fig. 8. Compared to our DARE model, TWIN-4E further decouples the behaviors and the target, meaning that the same category or item has two totally independent embeddings as behavior and target. This is strongly against two prior knowledges in recommendation system. 1. The correlation of two behaviors is similar no matter which is target and which is behavior. 2. Behaviors with the same category should be more correlated, which is natural in DARE since a vector’s dot product with itself tends to be bigger.

### v2

@@ Abstract
Lifelong user behavior sequences are crucial for capturing user interests and predicting user responses in modern recommendation systems. A two-stage paradigm is typically adopted to handle these long sequences: a subset of relevant behaviors is first searched from the original long sequences via an attention mechanism in the first stage and then aggregated with the target item to construct a discriminative representation for prediction in the second stage. In this work, we identify and characterize, for the first time, a neglected deficiency in existing long-sequence recommendation models: a single set of embeddings struggles with learning both attention and representation, leading to interference between these two processes. Initial attempts to address this issue with some common methods (e.g., linear projections—a technique borrowed from language processing) proved ineffective, shedding light on the unique challenges of recommendation models. To overcome this, we propose the Decoupled Attention and Representation Embeddings (DARE) model, where two distinct embedding tables are initialized and learned separately to fully decouple attention and representation. Extensive experiments and analysis demonstrate that DARE provides more accurate searches of correlated behaviors and outperforms baselines with AUC gains up to 9‰ on public datasets and notable improvements on Tencent’s advertising platform. Furthermore, decoupling embedding spaces allows us to reduce the attention embedding dimension and accelerate the search procedure by 50% without significant performance impact, enabling more efficient, high-performance online serving. Code in PyTorch for experiments, including model analysis, is available at https://github.com/thuml/DARE.
@@ 1 Introduction
In online services, system response delays can severely disrupt the user experience, making efficient handling of long sequences within a limited time crucial. A general paradigm employs a two-stage process (Pi et al., 2020): search (a.k.a. General Search Unit) and sequence modeling (a.k.a. Exact Search Unit). This method relies on two core modules: the attention module [^1] In this paper, “attention” refers to attention scores—the softmax output that weights each behavior., which measures the target-behavior correlation, and the representation module, which generates discriminative representations of behaviors. The search stage uses the attention module to retrieve top-k relevant behaviors, constructing a shorter sub-sequence from the original long behavior sequence [^2] The search stage can also be “hard” selecting behaviors by category, but we focus on soft search based on learned correlations for better user interest modeling.. The sequence modeling stage relies on both modules to predict user responses by aggregating behavior representations in the sub-sequence based on their attention, thus extracting a discriminative representation. Existing works widely adopt this paradigm (Pi et al., 2020; Chang et al., 2023; Si et al., 2024).
@@
Attention is critical in the long-sequence recommendation, as it not only models the importance of each behavior for sequence modeling but, more importantly, determines which behaviors are selected in the search stage. However, in most existing works, the attention and representation modules share the same embeddings despite serving distinct functions—one learning correlation scores, the other learning discriminative representations. We analyze these two modules, for the first time, in the perspective of Multi-Task Learning (MTL). (Caruana, 1997). Adopting gradient analysis commonly used in MTL (Yu et al., 2020; Liu et al., 2021), we reveal that, unfortunately, gradients of these shared embeddings are dominated by representation, and more concerning, gradient directions from two modules tend to conflict with each other. Domination and conflict of gradients are two typical phenomena of interference between tasks, influencing the model’s performance on both tasks. Our experimental results are consistent with the theoretical insight: attention fails to capture behavior importance accurately, causing key behaviors to be mistakenly filtered out during the search stage (as shown in Sec. 4.3). Furthermore, gradient conflicts also degrade the discriminability of the representations (as shown in Sec. 4.4).
@@
Inspired by the use of separate query, key (for attention), and value (for representation) projection matrices in the original self-attention mechanism (Vaswani et al., 2017), we experimented with attention- and representation-specific projections in recommendation models, aiming to resolve conflicts between these two modules. However, this approach did not yield positive results. We also tried three other kinds of candidate methods, but unfortunately, none of them worked effectively. Through insightful empirical analysis, we hypothesize that the failure is due to the significantly lower capacity (i.e., fewer parameters) of the projection matrices in recommendation models compared to those in natural language processing (NLP). This limitation is difficult to overcome, as it stems from the low embedding dimension imposed by interaction collapse theory (Guo et al., 2024).
@@
To address these issues, we propose the Decoupled Attention and Representation Embeddings (DARE) model, which completely decouples these two modules at the embedding level by using two independent embedding tables—one for attention and the other for representation. This decoupling allows us to fully optimize attention to capture correlation and representation to enhance discriminability. Furthermore, by separating the embeddings, we can accelerate the search stage by 50% by reducing the attention embedding dimension to half, with minimal impact on performance. On the public Taobao and Tmall long-sequence datasets, DARE outperforms the state-of-the-art TWIN model across all embedding dimensions, achieving AUC improvements of up to 9‰. Online evaluation on Tencent’s advertising platform, one of the world’s largest platforms, achieves a 1.47% lift in GMV (Gross Merchandise Value). Our contribution can be summarized as follows:
  • We identify the issue of interference between attention and representation learning in existing long-sequence recommendation models and demonstrate that common methods (e.g., linear projections borrowed from NLP) fail to decouple these two modules effectively.
  • We propose the DARE model, which uses module-specific embeddings to fully decouple attention and representation. Our comprehensive analysis shows that our model significantly improves attention accuracy and representation discriminability.
@@ 2 An In-Depth Analysis into Attention and Representation
In this section, we first review the general formulation for long-sequence recommendation. Then, we analyze the training of shared embeddings, highlighting the domination and conflict of gradients from the attention and representation modules. Finally, we explore why straightforward approaches (e.g., using module-specific projection matrices) fail to address the issue.
@@ 2.2 Gradient Analysis of Domination and Conflict
The attention and representation modules can be seen as two tasks: the former focuses on learning correlation scores for behaviors, while the latter focuses on learning discriminative (i.e., separable) representations in a high-dimensional space. However, current methods use a shared embedding for both tasks, which may cause a similar phenomenon to “task conflict” in Multi-Task Learning (MTL) (Yu et al., 2020; Liu et al., 2021) and prevent either from being fully achieved. To validate this assumption, we analyze the gradients from both modules on the shared embeddings.
@@
Following the methods in MTL, we empirically observe the gradients back propagated to the embeddings from the attention and representation modules. Comparing their gradient norms, we find that gradients from the representation are five times larger, dominating those from attention, as demonstrated in Fig. 2. Observing their gradient directions, we further find that in nearly two-thirds of cases, the cosine of the gradient angles is negative, indicating the conflict between them, as shown in Fig. 3. Domination and conflict are two typical phenomena of task interference, suggesting challenges in learning them well.
Figure 3: Cosine angles of gradients.
In summary, the attention module and representation modules optimize the embedding table towards different directions with varying intensities during training, causing attention to lose correlation accuracy and representation to lose its discriminability. Notably, due to domination, such influence is more severe to attention, as indicated by the poor learned correlation between categories in Sec. 4.3. While some commonly used techniques in MTL may ease the conflict, we tend to seek an optimized model structure that further resolves the conflict.
Finding 1. In sequential recommenders, gradients from the representation module tend to conflict with that from the attention module, and typically dominate the embedding gradients.
@@
Figure 4: Illustration and evaluation for adopting linear projections. (a-b) The attention module in the original TWIN and after adopting linear projections. (c) Performance of TWIN variants. Adopting linear projections causes an AUC drop of nearly 2% on Taobao.
@@ 2.3 Recommendation Models Call for More Powerful Decoupling Methods
Normal decoupling methods fail to resolve conflicts.
To address such conflict, a straightforward approach is to use separate projections for attention and representation, mapping the original embeddings into two new decoupled spaces. This is adopted in the standard self-attention mechanism (Vaswani et al., 2017), which introduces query, key (for attention), and value projection matrices (for representation). Inspired by this, we propose a variant of TWIN that utilizes linear projections to decouple attention and representation modules, named TWIN (w/ proj.). The comparison with the original TWIN structure is shown in Fig. 4a and 4b. Surprisingly, linear projection, which works well in NLP, loses efficacy in recommendation systems, leading to negative performance impact, as shown in Tab. 4c. We also tried three kinds of other candidate methods (MLP-based projection, strengthening the capacity of linear projection, and gradient normalization), resulting in a total of eight models, but none of them resolved the conflict effectively. For the structure of these models and more details, refer to Appendix C.
@@
The failure of introducing projection matrices makes us wonder why it works well in NLP but not in recommendation. One possible reason is that the relative capacity of projection matrices regarding the token numbers in NLP is usually strong, e.g., with an embedding dimension of 4096 in LLaMA3.1 (Dubey et al., 2024), there are around 16 million parameters ($4096\times 4096=16,777,216$) in each projection matrix to map only 128,000 tokens in the vocabulary. To validate our hypothesis, we conduct a synthetic experiment in NLP using nanoGPT (Andrej, ) with the Shakespeare dataset. In particular, we decrease its embedding dimension from 128 to 2 and check the performance gap between the two models with/without projection matrices. As shown in Fig. 5, we observe that when the matrix has enough capacity, i.e., the embedding dimension is larger than 16, projection leads to significantly less loss. However, when the matrix capacity is further reduced, the gap vanishes. Our experiment indicates that using projection matrices only works with enough capacity.
@@
In contrast, due to the interaction collapse theory (Guo et al., 2024), the embedding dimension in recommendation is usually no larger than 200, leading to only up to $40000$ parameters for each matrix to map millions to billions of IDs. Therefore, the projection matrices in recommendation never get enough capacity, making them unable to decouple attention and representation. In this case, other normal decoupling methods mentioned in Appendix C also suffer from weak capacity.
Finding 2. Normal methods like linear projection fail to decouple attention and representation in sequential recommendation models due to limited capacity caused by embedding dimension.
@@ 3 DARE: Decoupled Attention and Representation Embeddings
With all eight normal decoupling models shown in Appendix C failed, based on our analysis, we seek methods with enough capacity, hoping to completely resolve the conflict. To this end, we propose to decouple these two modules at the embedding level. That is, we employ two embedding tables, one for attention ($\bm{E}^{\text{Att}}$) and another for representation ($\bm{E}^{\text{Repr}}$). With gradient back propagated to different embedding tables, our method has the potential to fully resolve the gradient domination and conflict between these two modules. We introduce our model specifically in this section and demonstrate its advantage by experiments in the next section.
@@ 3.1 Attention Embedding
Attention measures the correlation between history behaviors and the target (Zhou et al., 2018). Following the common practice, we use the scaled dot-product function (Vaswani et al., 2017). Mathematically, the $i$-th history behavior $i$ and target $t$, are embedded into $\bm{e}_{i}^{\text{Att}},\bm{v}_{t}^{\text{Att}}\sim\bm{E}^{\text{Att}}$, where $\bm{E}^{\text{Att}}$ is the attention embedding table. After retrieval $1,2,\dots,K=\operatorname{Top-K}(\langle\bm{e}_{i},\bm{v}_{t}\rangle,i\in\left[1,N\right])$ their weight $w_{i}$ is formalized as:
@@
Figure 6: Architecture of the proposed DARE model. One embedding is responsible for attention, learning the correlation between the target and history behaviors, while another embedding is responsible for representation, learning discriminative representations for prediction. Decoupling these two embeddings allows us to resolve the conflict between the two modules.
@@ 3.2 Representation Embedding
In the representation part, another embedding table $\bm{E}^{\text{Repr}}$ is used, where $i$ and $t$ is embedded into $\bm{e}_{i}^{\text{Repr}},\bm{v}_{t}^{\text{Repr}}\sim\bm{E}^{\text{Repr}}$. Most existing methods multiply the attention weight with the representation of each retrieved behavior and then concatenate it with the embedding of the target as the input of Multi-Layer Perceptron (MLP): $[\sum_{i}w_{i}\bm{e}_{i},\bm{v}_{t}]$. However, it has been proved that MLP struggles to effectively learn explicit interactions (Rendle et al., 2020; Zhai et al., 2023). To enhance the discriminability, following TIN (Zhou et al., 2024), we adopt the target-aware representation $\bm{e}_{i}^{\text{Repr}}\odot\bm{v}_{t}^{\text{Repr}}$, denoted as TR in our following paper (refer to Sec. 4.4 for empirical evaluation of discriminability).
The overall structure of our model is shown in Fig. 6. Formally, user history $h$ is compressed into: $\bm{h}=\sum_{i=1}^{K}w_{i}\cdot(\bm{e}_{i}^{\text{Repr}}\odot\bm{v}_{t}^{\text{Repr}}).$
@@ 3.3 Inference Acceleration
By decoupling the attention and representation embedding tables, the dimension of attention embeddings $\bm{E}^{\text{Att}}$ and the dimension of representation embeddings $\bm{E}^{\text{Repr}}$ have more flexibility. In particular, we can reduce $\bm{E}^{\text{Att}}$ while keeping $\bm{E}^{\text{Repr}}$ to accelerate the searching over the original long sequence whilst not affecting the model’s performance. Empirical experiments in Sec. 4.5 show that our model has the potential to speed up searching by 50% with quite little influence on performance and even by 75% with an acceptable performance loss.
@@ 3.4 Discussion
Considering the superiority of decoupling the attention and representation embeddings, one may naturally raise an idea: we can further decouple the embeddings of history and target within the attention (and representation) module, i.e. forming a TWIN with 4 Embeddings method, or TWIN-4E in short, consisting of attention-history (named keys in NLP) $\bm{e}_{i}^{\text{Att}}\in\bm{E}^{\text{Att-h}}$, attention-target (named querys in NLP) $\bm{v}_{t}^{\text{Att}}\in\bm{E}^{\text{Att-t}}$, representation-history (named values in NLP) $\bm{e}_{i}^{\text{Repr}}\in\bm{E}^{\text{Repr-h}}$ and representation-target $\bm{v}_{t}^{\text{Repr}}\in\bm{E}^{\text{Repr-t}}$. The structure of TWIN-4E is shown in Fig. 7. Compared to our DARE model, TWIN-4E further decouples the behaviors and the target, meaning that the same category or item has two totally independent embeddings as behavior and target. This is strongly against two prior knowledge in recommendation system. 1. The correlation of two behaviors is similar no matter which is the target and which is from history. 2. Behaviors with the same category should be more correlated, which is natural in DARE since a vector’s dot product with itself tends to be bigger.